OpenStudy (anonymous):

Help solve please medal! Simplify 3x+10/ x+1 + x+3/x+4 -6x+27/(x+1)(x+4)

OpenStudy (anonymous):

@radar @sammixboo @MrNood @puppylife101

OpenStudy (anonymous):

@Yungwisdom2000 @Compassionate

OpenStudy (jtvatsim):

I'm going to rewrite this to make sure we are looking at the same thing. :)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (jtvatsim):

\[\frac{3x+10}{ x+1} + \frac{ x+3}{x+4} - \frac{6x+27}{(x+1)(x+4)}\] Is that the right thing?

OpenStudy (anonymous):

correct

OpenStudy (jtvatsim):

It's a bit scary looking. :) Any thoughts on what you might do first?

OpenStudy (anonymous):

we need a common denominator right :)

OpenStudy (jtvatsim):

That would be REALLY nice. Yes, let's try that.

OpenStudy (jtvatsim):

What might be the common denominator we are trying to make?

OpenStudy (anonymous):

(x+1) (x+4) right

OpenStudy (jtvatsim):

Seems like a good choice. That appears to be the most complicated thing in our expression. So far, so good.

OpenStudy (jtvatsim):

So, let's look at the first fraction. What needs to be done here?

OpenStudy (anonymous):

3x+10 (x+4) / x+1 (x+4) + (x+3) (x+1) / (x+4)(x+4) right

OpenStudy (jtvatsim):

Yes, on the second fraction you mean (x + 1) on the bottom, but I get it. :)

OpenStudy (jtvatsim):

Great! So here's what we have so far.

OpenStudy (jtvatsim):

\[\frac{(3x+10) (x + 4)}{(x+1) (x + 4)}+\frac{(x+3)(x+1)}{(x+1)(x+4)}−\frac{6x+27}{(x+1)(x+4)}\]

OpenStudy (jtvatsim):

Wow, that's a lot of stuff. Now what?

OpenStudy (anonymous):

cancel out

OpenStudy (jtvatsim):

Well, if we "cancel out" that might put us right back where we started right? Unless you mean distribute...

OpenStudy (anonymous):

i'm sorry i meant distribute lol

OpenStudy (jtvatsim):

No worries. :) Let's try to do that. What do you get. (Might take a minute) :)

OpenStudy (anonymous):

im kinda lost can you draw it out

OpenStudy (anonymous):

from where we left off

OpenStudy (jtvatsim):

Sure! I think we were here right? \[\frac{(3x+10)(x+4)}{(x+1)(x+4)}+\frac{(x+3)(x+1)}{(x+1)(x+4)}−\frac{6x+27}{(x+1)(x+4)}\]

OpenStudy (anonymous):

yes ok, so we distribute

OpenStudy (jtvatsim):

Yes, and what do you get after that?

OpenStudy (anonymous):

i got 4/ (x+1) (x+4)

OpenStudy (jtvatsim):

4 for which fraction?

OpenStudy (jtvatsim):

As the final answer?

OpenStudy (anonymous):

yes the final

OpenStudy (anonymous):

i think i did something wrong

OpenStudy (jtvatsim):

Oh, ok, let me double check that. :)

OpenStudy (anonymous):

yes please check

OpenStudy (jtvatsim):

So, for the first fraction: \[\frac{(3x+10)(x+4)}{(x+1)(x+4)}=\frac{3x^2 + 12x + 10 x + 40}{(x+1)(x+4)}\]

OpenStudy (jtvatsim):

For the second fraction: \[\frac{(x+3)(x+1)}{(x+1)(x+4)} = \frac{x^2 + 3x + x + 3}{(x+1)(x+4)}\]

OpenStudy (jtvatsim):

For the third fraction (watch out for the tricky - sign!): \[-\frac{6x+27}{(x+1)(x+4)}=\frac{-6x-27}{(x+1)(x+4)}\]

OpenStudy (anonymous):

ok i got 1/ (x+4)(x+1) for final

OpenStudy (jtvatsim):

Checking again... :)

OpenStudy (jtvatsim):

I get: (3x^2 + x^2 + 22x + 4x - 6x + 40 + 3 - 27)/(x+1)(x+4) so (4x^2 + 20x + 16)/(x+1)(x+4) so (2x + 2)(2x + 8)/(x+1)(x+4) then 2(x+1)*2(x+4)/(x+1)(x+4) Oh wow! 4 is the final answer. I think. :)

OpenStudy (anonymous):

ok i see, now you factor

OpenStudy (anonymous):

hey do you mind checking this one for me

OpenStudy (anonymous):

x/3 div 2/x+5

OpenStudy (anonymous):

i got X(x+5) /6